OpenStudy (anonymous):

Does the table below represent a function? Car wash grand opening time, X water,Y (minutes) (gallons) 0 440 1 428 2 416 3 404 A. yes B. no @iTrinity @iGreen @ASAAD123 @animal_lover36

OpenStudy (anonymous):

A.Yes

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

@QueenCandy

OpenStudy (anonymous):

A.yes \[m=\frac{ 428-440 }{ 1-0}=-12\] y-428=-12(x-1) y=-12x+440